OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence.a16 = 21, a17 = -1

OpenStudy (***[isuru]***):

the general equation for the nt term of an arithmetic progression is \[T _{n} = a + ( n -1 ) d\] where " Tn" is the n th term " a" is the first term "d" is the common difference... so... from ur data a 16 = 21 a 17 = -1 apply this data to the above formula.. for a16 Tn = a + (n-1)d a16 = a + (16 -1)d 21 = a + 15d ------(1) for a17 Tn = a + (n-1)d a17 = a + (17 -1)d -1 = a + 16d ----(2) now (1) - (2) 21- (-1) = a + 15d - ( a + 16d) 22 = a + 15d -a -16d 22 = -d d = -22 now substitute d =(-22) in (1) 21 = a + [15x(-22)] 21= a -330 a = 330 +21 a = 351 so... now u know the values a =351 d = -22 so.. the nth term is... \[T _{n} = a + ( n -1 ) d\] \[T _{n} = 351 - ( n -1 ) 22\] hope this will help ya!!!

OpenStudy (anonymous):

Is this the same as your answer an = 351 - 22(n - 1)

OpenStudy (***[isuru]***):

yep!!

OpenStudy (anonymous):

Thank You !